Based on a sample of n = 20, the least-squares method was used to develop the following prediction line: Yi = 5 + 3Xi

In addition,

SYX = 1.0

\bar{X} = 2

\bg_white \fn_jvn \sum_{i = 1}^{n} (X_{I}-\bar{X})^2 = 20

a) Construct a 95% confidence interval estimate of the population mean response for X = 2.

b) Construct a 95% prediction interval of an individual response for x = 2.
Answer #1
n= 20
bo= 5
b1= 3.0000
sxy =√MSE= 1.0000
Sxx=(n-1)sx^2= 20
x̅ = 2

a)

x   = 2
predcited value at X=2: 11.0000
std error confidence interval= s*√(1/n+(x0-x̅)2/Sxx) = 0.2236
for 95 % CI value of t= 2.1010
margin of error E=t*std error                            = 0.4698
lower confidence bound=sample mean-margin of error = 10.5302
Upper confidence bound=sample mean+margin of error= 11.4698

95% confidence interval =(10.5302 , 11.4698) (please use decimals as required in question)

b)

x0             = 2
predcited value at X=2: 11.00
std error prediction interval= s*√(1+1/n+(x0-x̅)2/Sxx) = 1.0247
for 95 % CI value of t= 2.1010
margin of error E=t*std error                            = 2.15
lower prediction bound=sample mean-margin of error = 8.8471
Upper prediction bound=sample mean+margin of error= 13.1529

95% prediction interval =(8.8471 , 13.1529)